# PRISM: Prompt-Refined In-Context System Modelling for Financial Retrieval

### v1

@@ A.2 Dynamics of In-Context Learning and Prompt Versioning
ICL was introduced only at the final prompt stage to maintain a clear, experiment-driven workflow. We first focused on stabilizing the prompt design, since it governs the model’s reasoning structure and output constraints. Once a stable and effective prompt was established, we introduce ICL as an extension to assess how contextual examples could further enhance model performance. The ICL design evolved progressively, starting with 5 examples to test baseline adaptability, then expanding to 10 and 15 examples to assess the model’s ability to leverage larger context windows. This incremental setup allowed us to observe how additional examples influence reasoning stability and retrieval accuracy. By introducing ICL only after the prompt was fully optimized, we ensured that any observed improvements could be attributed directly to the contextual learning mechanism rather than prompt variation, maintaining consistency and preserving interpretability across experiments.
A.3 Implementation Details
The exact versions of the four foundation models used in our experiments are provided below:
1. GPT-4o-mini: gpt-4o-mini-2024-07-18
2. GPT-4.1: gpt-4.1-2025-04-14
3. GPT-5-mini: gpt-5-mini-2025-08-07
4. GPT-5: gpt-5-2025-08-07
The retrieval pipeline was implemented using a FAISS vector store with two OpenAI’s embedding backbones: text-embedding-3-small v1 (TE3-S) and text-embedding-3-large (TE3-L). Multi-agent workflows were constructed with LangGraph (v1.0.3), and all models were accessed through the OpenAI Python SDK (v2.3.0).
A.3.1 Model Provider Selection
@@
All experiments were conducted exclusively using OpenAI models. This decision was driven primarily by practical and infrastructural considerations as most companies in our region operate within the Microsoft Azure ecosystem, which provides direct and optimized access to OpenAI suite of models. As a startup, we leveraged limited Azure AI credits and native integration to efficiently run large-scale experiments with minimal development and deployment overhead. Beyond operational accessibility, OpenAI’s remain among the most benchmarked closed-source LLMs on financial tasks bigeard2025finance, offering strong reliability and reproducibility. Therefore, we believe using OpenAI as the sole model provider ensures a stable experimental environment and a credible baseline for future comparisons with alternative LLMs.
A.3.2 LangGraph Selection
@@
LangGraph was selected to manage multi-agentic workflows due to its modular design, strong community support, and active maintenance. As a widely adopted framework, it provides a stable and extensible foundation for orchestrating complex agent interactions, allowing developers to define and control the flow of information between agents with precision. Its high degree of customizability enables flexible integration with diverse tools, APIs, and models, making it adaptable to various use cases. Besides, it emphasizes on maintainability and observability that supports both rapid prototyping during development and scalable deployment in production. Therefore, these factors make LangGraph an ideal choice for orchestrating complex MAS financial reasoning pipelines while ensuring long-term sustainability in both research and applied settings.
A.3.3 Prompt Selection in Agentic Workflow.
@@
We fixed the prompt version to $P_{1}$ for both chunk- and document-ranking tasks in all agentic workflow experiments to ensure controlled and interpretable results. This allowed us to isolate the effects of architectural modifications, ensuring that any performance differences could be attributed to changes in agent coordination rather than confounded by prompt variation. Furthermore, system prompts in a multi-agent setup serve distinct roles across agents, and later prompt versions contain more complex instructions that do not generalize well across heterogeneous agent responsibilities. For example, in Run 32, performance deteriorated when applying the $P_{4}$ prompt to the document-ranking task, as the increased reasoning complexity introduced confusion among agents. Similarly, attempting to use $P_{4}$ for chunk-ranking tasks in the hybrid workflow revealed practical constraints as the expanded prompt frequently exceeded token limits due to the large chunk sizes. For these reasons, we maintained a consistent and lightweight prompt configuration while experimenting with different agentic architectures, ensuring comparability, scalability, and clarity in performance attribution.
A.4 Quantitative Results
Figure 6: Top 10 teams ranked by private subset scores.
Quantitative results of top 10 teams evaluated on private subset are illustrated in Figure 6. The results also show that the top 3 methods have a small gap between each other and have bigger gaps compared to other methods.
A.5 Feasibility Analysis
A.5.1 Cost Analysis

### v2

@@ A.5 Dynamics of In-Context Learning and Prompt Versioning
ICL was introduced only at the final prompt stage to maintain a clear, experiment-driven workflow. We first focused on stabilizing the prompt design, since it governs the model’s reasoning structure and output constraints. Once a stable and effective prompt was established, we introduced ICL as an extension to assess how contextual examples could further enhance model performance. The ICL design evolved progressively, starting with 5 examples to test baseline adaptability, then expanding to 10 and 15 examples to assess the model’s ability to leverage larger context windows. This incremental setup allowed us to observe how additional examples influence reasoning stability and retrieval accuracy. By introducing ICL only after the prompt was fully optimized, we ensured that any observed improvements could be attributed directly to the contextual learning mechanism rather than prompt variation, maintaining consistency and preserving interpretability across experiments.
A.6 In-Context Learning Retrieval Details
The FAISS store used for ICL exemplar retrieval is separate from the document/chunk corpus that the system ranks; the former provides learning examples while the latter contains the actual candidates to be ranked. This separation ensures that retrieved exemplars serve as reasoning scaffolds without contaminating the ranking candidates. The resulting prompt is dynamic, adapting to the semantic characteristics of each user query rather than relying on a static template. Empirically, we find that ICL performs best when the retrieved samples are semantically close to the target query, as closer exemplars provide more relevant reasoning patterns for the model to follow.
A.7 Multi-Agent System Architecture Details
This section provides comprehensive implementation details for the Multi-Agent System (MAS) modeling architectures used in PRISM, including the document ranking workflow and four chunk ranking variants ($A_{1}$$A_{4}$). All architectures are illustrated in Figure 7.
A.7.1 Agent Definitions
We adopt a state-graph framework implemented using LangGraph, where each agent is an LLM instance with a specialized system prompt defining its role and scoring criteria. Document ranking employs a fixed configuration: a Question Analyzer that extracts key entities and intent from the query, followed by five Document Expert agents (one per SEC filing type: 10-K, 10-Q, 8-K, DEF 14A, and Earnings Transcripts) that score documents based on domain-specific relevance criteria. For chunk ranking, we define specialized agents including: Financial Analyst (evaluates quantitative metrics and financial data), Risk Analyst (assesses operational and regulatory risks), Evidence Extractor (identifies concrete supporting facts), Contextual Reasoner (evaluates explanatory content), and filtering agents such as Quick Filter and Noise Remover.
A.7.2 Interaction Process
Agents communicate through a shared state graph where each node represents an agent and edges define information flow. In a typical workflow: (1) input chunks are passed to filtering agents that assign preliminary relevance scores; (2) filtered candidates flow to scoring agents that independently evaluate chunks on a 1–10 scale based on their specialized criteria; (3) scores are aggregated through weighted averaging or voting to produce final rankings. The graph topology determines whether agents operate in parallel (simultaneous scoring) or sequentially (pipeline filtering). All agent prompts are provided in Table LABEL:tab:agentic-chunk-prompts.
A.7.3 State Management and Information Flow
All MAS architectures maintain a shared state graph implemented through LangGraph’s StateGraph abstraction. Information flows through the graph according to predefined edges connecting agent nodes. Parallel agents (e.g., scoring agents in $A_{1}$) execute simultaneously without inter-agent communication, reducing coordination complexity. Sequential pipelines ($A_{2}$, $A_{3}$) enforce strict ordering where downstream agents only receive filtered candidates from upstream stages, creating potential bottlenecks if early filtering is overly aggressive. The state contains the input data, intermediate outputs, and aggregated results.
A.7.4 Document Ranking Workflow
The document ranking workflow (Figure 7(a)) follows a two-stage architecture designed to handle heterogeneous financial document types with specialized domain expertise. The workflow begins with a Question Analyzer agent that processes the user query to extract key entities and classify the query relevance to the corresponding agents’ expertise. This structured analysis is then distributed to five parallel Document Expert agents, each specialized in a specific SEC filing type:
  • 10-K Expert: Evaluates annual reports, focusing on comprehensive financial statements, business operations, and risk factors disclosed in Form 10-K filings.
  • 10-Q Expert: Assesses quarterly reports, prioritizing interim financial performance, management discussion and analysis (MD&A), and quarterly updates.
  • 8-K Expert: Analyzes current reports for material events, corporate announcements, and time-sensitive disclosures.
  • DEF14A Expert: Reviews proxy statements for governance information, executive compensation, and shareholder proposals.
  • Earnings Transcript Expert: Evaluates earnings call transcripts for management commentary, forward guidance, and analyst Q&A insights.
Each document expert scores documents on a 0–4 scale based on relevance to the analyzed query intent. Scores are then aggregated through weighted averaging, where weights are determined by document type relevance to the query category provided by the question analyzer. This parallel scoring approach ensures domain-appropriate evaluation while maintaining computational efficiency, as document experts operate simultaneously without sequential dependencies.
A.7.5 Chunk Ranking Workflow Variants
The chunk ranking workflows (Figures 7(b)7(e)) represent four architectural variants with progressively simplified designs, each exploring different trade-offs between reasoning depth, error propagation, and computational cost.
$A_{1}$: Direct Scoring Ensemble (Figure 7(b)) implements a single-stage parallel scoring approach with four domain-expert agents operating simultaneously:
  • CEO Agent: Evaluates strategic relevance and high-level business implications.
  • Financial Analyst: Assesses quantitative metrics, financial ratios, and numerical evidence.
  • Operations Manager: Focuses on operational details, business processes, and execution aspects.
  • Risk Analyst: Identifies risk factors, uncertainties, and potential adverse impacts.
Each agent independently scores all candidate chunks on a 1–10 scale according to its specialized criteria. Final rankings are determined through simple averaging of the four scores. This architecture minimizes pipeline depth to reduce error accumulation and analyzes chunks from different perspectives.
$A_{2}$: Three-Stage Deep Pipeline (Figure 7(c)) implements the most complex workflow with three sequential stages designed for sequential filtering and specialized evaluation:
  1. 1.
    Noise Removal Stage: A Noise Remover agent filters out clearly irrelevant chunks (boilerplate text, legal disclaimers, table fragments) to reduce downstream computational burden.
  2. 2.
    Candidate Selection Stage: A Candidate Selector agent performs secondary filtering, retaining only chunks with potential relevance based on keyword matching and semantic proximity.
  3. 3.
    Ensemble Scoring Stage: Four specialized agents score the filtered candidates:
    • Relevance Scorer: Direct query-chunk semantic alignment.
    • Contextual Reasoner: Evaluates explanatory content and contextual completeness.
    • Evidence Extractor: Identifies concrete supporting facts and data points.
    • Diversity Agent: Ensures ranking diversity to avoid redundant information.
While this deep architecture aims to progressively refine candidates, our experiments reveal it suffers from over-filtering and cascading errors, where early-stage false negatives cannot be recovered in later stages, leading to degraded recall performance.
$A_{3}$: Two-Stage Balanced Pipeline (Figure 7(d)) simplifies $A_{2}$ by reducing to two stages while maintaining filtering benefits:
  1. 1.
    Quick Filtering Stage: A Quick Filter agent performs lightweight relevance assessment, applying looser thresholds than $A_{2}$’s Noise Remover to preserve recall.
  2. 2.
    Specialized Scoring Stage: Three focused agents evaluate filtered chunks:
    • Relevance Scorer: Query-chunk alignment assessment.
    • Contextual Reasoner: Contextual adequacy and explanatory quality.
    • Evidence Extractor: Concrete evidence identification.
This architecture balances pipeline depth with stability, reducing error propagation risks while maintaining some computational efficiency gains from preliminary filtering.
$A_{4}$: Minimal Two-Agent System (Figure 7(e)) represents the simplest viable MAS configuration with direct parallel scoring by two complementary agents:
  • Financial Analyst: Focuses on quantitative evidence and financial metrics.
  • Risk Analyst: Evaluates qualitative risk factors and uncertainties.
By eliminating filtering stages and reducing agent count, it minimizes interaction complexity and error propagation pathways. This minimalist design performed comparably to or better than deeper architectures in our experiments, as reduced error propagation outweighs ensemble benefits.
(a) Document ranking workflow
(b) $A_{1}$: Direct scoring ensemble
(c) $A_{2}$: Three-stage deep pipeline
(d) $A_{3}$: Two-stage balanced pipeline
(e) $A_{4}$: Minimal two-agent system
Figure 7: Multi-Agent System architectures for PRISM. (a) Document ranking workflow. (b)–(e) Chunk ranking variants.
A.7.6 Prompt Engineering for Agent Roles
Each agent receives a specialized system prompt defining its role, evaluation criteria, and output format. For document ranking, expert prompts specify document-type-specific relevance factors. For chunk ranking, scoring agents receive detailed rubrics mapping scores 1–10 to specific evidence characteristics. Filtering agents use binary decision criteria with explicit instructions to prefer recall over precision. All agent prompts are provided in Table LABEL:tab:agentic-doc-prompts and Table LABEL:tab:agentic-chunk-prompts.
A.7.7 Aggregation and Ranking Logic
Score aggregation varies by architecture. Simple averaging is used in chunk ranking where all agents have equal weight. In document ranking, we implement weighted averaging where weights reflect agent reliability and query-dependent relevance. Final rankings are produced by sorting candidates by aggregated score in descending order, with ties broken by original retrieval rank.
A.8 Implementation Details
A.8.1 Model Provider Selection
@@
All experiments were conducted exclusively using OpenAI models. This decision was driven primarily by practical and infrastructural considerations as most companies in our region operate within the Microsoft Azure ecosystem, which provides direct and optimized access to OpenAI’s suite of models. As a startup, we leveraged limited Azure AI credits and native integration to efficiently run large-scale experiments with minimal development and deployment overhead. Beyond operational accessibility, OpenAI’s models remain among the most benchmarked closed-source LLMs on financial tasks (Bigeard et al., 2025), offering strong reliability and reproducibility. Therefore, we believe using OpenAI as the sole model provider ensures a stable experimental environment and a credible baseline for future comparisons with alternative LLMs. The exact versions of the four foundation models used in our experiments are provided below:
  1. 1.
    GPT-4o-mini: gpt-4o-mini-2024-07-18
  2. 2.
    GPT-4.1: gpt-4.1-2025-04-14
  3. 3.
    GPT-5-mini: gpt-5-mini-2025-08-07
  4. 4.
    GPT-5: gpt-5-2025-08-07
The retrieval pipeline was implemented using a FAISS vector store with two OpenAI embedding backbones: text-embedding-3-small v1 (TE3-S) and text-embedding-3-large (TE3-L). Multi-agent workflows were constructed with LangGraph (v1.0.3), and all models were accessed through the OpenAI Python SDK (v2.3.0).
A.8.2 LangGraph Selection
@@
LangGraph was selected to manage multi-agentic workflows due to its modular design, strong community support, and active maintenance. As a widely adopted framework, it provides a stable and extensible foundation for orchestrating complex agent interactions, allowing developers to define and control the flow of information between agents with precision. Its high degree of customizability enables flexible integration with diverse tools, APIs, and models, making it adaptable to various use cases. Additionally, its emphasis on maintainability and observability supports both rapid prototyping during development and scalable deployment in production. These factors make LangGraph a suitable choice for orchestrating complex MAS financial reasoning pipelines while ensuring long-term sustainability in both research and applied settings.
A.8.3 Prompt Selection in Agentic Workflow
@@
We fixed the prompt version to $P_{1}$ for both chunk- and document-ranking tasks in all agentic workflow experiments to ensure controlled and interpretable results. This allowed us to isolate the effects of architectural modifications, ensuring that any performance differences could be attributed to changes in agent coordination rather than confounded by prompt variation. Furthermore, system prompts in a multi-agent setup serve distinct roles across agents, and later prompt versions contain more complex instructions that do not generalize well across heterogeneous agent responsibilities. For example, in Run 32, performance deteriorated when applying the $P_{4}$ prompt to the document-ranking task, as the increased reasoning complexity introduced confusion among agents. Similarly, attempting to use $P_{4}$ for chunk-ranking tasks in the hybrid workflow revealed practical constraints as the expanded prompt frequently exceeded token limits due to the large chunk sizes. For these reasons, we maintained a consistent and lightweight prompt configuration while experimenting with different agentic architectures, ensuring comparability, scalability, and clarity in performance attribution. It is worth noting that this setup introduces a prompt asymmetry: agentic experiments use $P_{1}$ throughout, while the best non-agentic result uses $P_{4}$. We discuss this design choice and its implications further in Appendix A.13.3.
A.9 Quantitative Results
A.9.1 FinAgentBench
Table 9: FinAgentBench leaderboard results (Private NDCG@5 scores).
Rank Team NDCG@5 Score (Private) Open Source Fine-tuning
1 memex 0.72497 N/A Yes
2 Yuzhen Hu 0.72328 N/A Yes
[…]
4 rlawnsxo 0.70601 N/A N/A
5 currio 0.70195 N/A N/A
6 kaggleslasher 0.69985 N/A N/A
7 Gautam Jajoo 0.69774 N/A N/A
8 RagTag 0.69173 N/A N/A
9 ART YOUNG 0.69162 N/A N/A
10 liangjiang chen 0.69081 N/A N/A
[…]
Table 9 reports the final private-subset leaderboard rankings from the FinAgentBench competition. We include this table for transparency and contextual framing of our competition placement, since the evaluation environment is no longer available. The top three methods show a small performance gap relative to the remaining teams. The memex team’s solution employed a pool-of-experts of five different LLMs for pre-filtering and 65 agents for chunk retrieval and reranking, with an embedding model fine-tuned for 30 epochs on FinAgentBench training data. The second-ranked Yuzhen Hu team’s solution used three MiniLM models fine-tuned separately for document retrieval, chunk retrieval, and cross-encoder reranking. In contrast, PRISM is training-free and relies exclusively on off-the-shelf LLMs with engineered prompts and lightweight multi-agent coordination. The performance gap between PRISM and the top two fine-tuned approaches ($\Delta\approx 0.005$$0.007$ NDCG@5 on the private subset) is modest and reflects the expected trade-off of forgoing task-specific fine-tuning.
Chunk Ranking Analysis.
While Table 3 reports combined NDCG@5 scores, our detailed analysis reveals that chunk ranking is substantially more challenging than document ranking. Chunk ranking accounts for the majority of latency (131s vs. 10s for documents in Run 19) and token usage (approximately 95K prompt tokens vs. 3K for documents). The performance gap between agentic and non-agentic configurations is most pronounced at the chunk level: MAS architectures consistently underperformed non-agentic approaches for chunk ranking due to error propagation across sequential agent stages. Our best chunk ranking results were achieved with the $P_{4}$ prompt and GPT-5 without ICL augmentation, suggesting that for fine-grained chunk discrimination, a well-designed single-model approach outperforms distributed multi-agent reasoning. This finding has practical implications: practitioners should prioritize prompt optimization and model selection for chunk ranking, reserving MAS for document-level tasks where coordination overhead is better tolerated.
A.9.2 FIQA-2018
Table LABEL:tab:full_fiqa_results presents comprehensive retrieval performance metrics on FIQA-2018, revealing PRISM’s substantial superiority over traditional baselines (Table 10) across all configurations. Most notably, PRISM consistently surpasses the best-performing baseline (BM25+CE) by large margins, with even the lowest-performing PRISM variant achieving better results.
Examining the prompt design progression reveals nuanced performance patterns. $P_{1}$, employing a straightforward prompting strategy, establishes a strong baseline with GPT-5 achieving 0.6130 NDCG@10 and 0.8045 Recall@100. $P_{2}$, which integrates key contextual information, demonstrates comparable performance of 0.6095 NDCG@10 and 0.8094 Recall@100 with GPT-5, suggesting that additional context maintains effectiveness while potentially offering enhanced interpretability. $P_{3}$, utilizing Chain-of-Thought (CoT) reasoning, achieves the highest NDCG@10 of 0.6197, indicating that structured reasoning particularly benefits ranking precision at top positions. Interestingly, $P_{4}$, which combines ReAct and Tree-of-Thought (ToT) for more explicit multi-step reasoning, shows slightly lower performance of 0.6070 NDCG@10, suggesting that excessive prompt complexity may introduce overhead without proportional gains for this task.
The impact of ICL varies significantly by model capability, revealing important insights for practical deployment. ICL proves particularly beneficial for models with lesser reasoning capabilities, as evidenced by GPT-4.1’s consistent performance gains across all prompts (e.g., 23.5% improvement in NDCG@10 for $P_{1}$ configuration). This enhancement occurs because ICL provides concrete demonstrations of the task structure and expected reasoning patterns, effectively compensating for limited inherent reasoning capacity by offering explicit templates to follow. In contrast, GPT-5-mini and GPT-5 exhibit a more stable performance with marginal variations when ICL is applied, indicating their robust capacity to generalize from task instructions alone without requiring extensive demonstration examples unless it is a highly specialized task. This differential ICL sensitivity suggests that smaller or less capable models can be substantially enhanced through careful example selection, while more advanced models already possess sufficient task understanding capabilities.
Overall, all PRISM configurations maintain higher Recall@100 ($\geq 0.80$ for GPT-5) than the best-performing baseline, demonstrating strong retrieval coverage, while improvements in NDCG@10 highlight enhanced precision in ranking highly relevant documents at top positions, a critical capability for practical information retrieval systems.
Table 10: Traditional baseline performance on FIQA-2018.
Method Model NDCG@10 Recall@100 BM25 Baseline 0.2360 0.5390 DeepCT Sparse 0.1910 0.4890 SPARTA Sparse 0.1980 0.4460 docT5query Sparse 0.2910 0.5980 DPR Dense 0.1120 0.3420 ANCE Dense 0.2950 0.5810 TAS-B Dense 0.3000 0.5930 GenQ Dense 0.3080 0.6180 ColBERT Late-Interaction 0.3170 0.6030 BM25+CE Re-ranking 0.3470 0.5390
Table 11: PRISM performance on FIQA-2018.
Prompt Model ICL NDCG@10 Recall@100
$P_{1}$ GPT-4.1 N/A 0.4307 0.6024
$P_{1}$ GPT-5-mini N/A 0.5565 0.7268
$P_{1}$ GPT-5 N/A 0.6130 0.8045
$P_{1}$ GPT-4.1 ICL-5 0.5320 0.7247
$P_{1}$ GPT-5-mini ICL-5 0.5171 0.6412
$P_{1}$ GPT-5 ICL-5 0.6104 0.8083
$P_{2}$ GPT-4.1 N/A 0.4871 0.6711
$P_{2}$ GPT-5-mini N/A 0.5521 0.7094
$P_{2}$ GPT-5 N/A 0.6095 0.8094
$P_{2}$ GPT-4.1 ICL-5 0.4974 0.6863
$P_{2}$ GPT-5-mini ICL-5 0.4960 0.6142
$P_{2}$ GPT-5 ICL-5 0.6027 0.7941
$P_{3}$ GPT-4.1 N/A 0.4479 0.5948
$P_{3}$ GPT-5-mini N/A 0.5642 0.7374
$P_{3}$ GPT-5 N/A 0.6094 0.8136
$P_{3}$ GPT-4.1 ICL-5 0.4702 0.6471
$P_{3}$ GPT-5-mini ICL-5 0.5182 0.6506
$P_{3}$ GPT-5 ICL-5 0.6197 0.8206
$P_{4}$ GPT-4.1 N/A 0.4178 0.5739
$P_{4}$ GPT-5-mini N/A 0.5269 0.6812
$P_{4}$ GPT-5 N/A 0.6070 0.8195
$P_{4}$ GPT-4.1 ICL-5 0.4544 0.6212
$P_{4}$ GPT-5-mini ICL-5 0.4615 0.5610
$P_{4}$ GPT-5 ICL-5 0.6097 0.8021
A.9.3 FinanceBench
To assess PRISM’s versatility beyond reranking tasks, we extended our evaluation to FinanceBench, complementing our experiments on the reranking datasets FiQA and FinAgentBench. For a reliable and fair comparison, we replicated the best-performing baseline from the FinanceBench paper, implementing their LLM approach with similar models across multiple evaluation modes. Table LABEL:tab:full_financebench_results presents comprehensive results revealing PRISM’s effectiveness in question-answering tasks. Under the Oracle context conditions, PRISM achieved a near-perfect accuracy of 98%, substantially surpassing the baseline performance. More importantly, PRISM demonstrates consistent improvements across all evaluation modes and models. For instance, all PRISM variants ($P_{1}$$P_{4}$) achieve 97%–98% accuracy, representing improvement in an already high-performing regime.
An important observation emerges regarding prompt complexity and task alignment. While $P_{4}$ showed slightly lower performance in reranking tasks, it maintains competitive performance in QA settings, with GPT-5 Oracle achieving 97.33% accuracy (146/150). This suggests that optimal prompt design is indeed task-dependent, as we observed in the reranking experiments. Simpler prompts ($P_{1}$) often match or exceed more complex variants, indicating that the structured reasoning scaffolding beneficial for nuanced ranking decisions may be less critical for direct factual retrieval tasks inherent to question-answering.
The limited utility of ICL in FinanceBench can be attributed to task-specific characteristics unlike its substantial benefits in reranking tasks. FinanceBench queries require finding specific answers within individual documents rather than comparing multiple candidates, making similarity-based ICL selection less effective at retrieving truly relevant question-answer exemplars. This results in ICL examples that may misdirect rather than guide the model, explaining why non-ICL configurations often perform comparably or better. Nevertheless, despite the limited dataset size (150 questions) and the already high baseline performance of up to 96% accuracy, PRISM achieves measurable improvements, demonstrating its effectiveness across different task types beyond reranking. While the absolute gains are necessarily constrained in such a high-performing regime approaching the theoretical ceiling of 100%, PRISM’s consistent ability to push accuracy toward 98% validates its versatility as a general-purpose framework for LLM-based information retrieval tasks.
Table 12: Comprehensive evaluation results of different prompt versions and models across various evaluation modes on FinanceBench.
Prompt Model Eval Mode ICL Results (Out of 150)
Correct Incorrect Unable to Answer
Baselines
Baseline GPT-4.1 Single Store N/A 132 18 0
Baseline GPT-4.1 Shared Store N/A 124 26 0
Baseline GPT-4.1 In Context N/A 129 21 0
Baseline GPT-4.1 Oracle N/A 133 16 1
Baseline GPT-5-mini Single Store N/A 144 4 2
Baseline GPT-5-mini Shared Store N/A 143 7 0
Baseline GPT-5-mini In Context N/A 142 7 1
Baseline GPT-5-mini Oracle N/A 145 5 0
Baseline GPT-5 Single Store N/A 141 8 1
Baseline GPT-5 Shared Store N/A 143 7 0
Baseline GPT-5 In Context N/A 139 11 0
Baseline GPT-5 Oracle N/A 141 9 0
PRISM
$P_{1}$ GPT-4.1 Single Store ICL-9 124 26 0
$P_{1}$ GPT-4.1 Shared Store ICL-9 126 24 0
$P_{1}$ GPT-4.1 In Context ICL-9 124 26 0
$P_{1}$ GPT-4.1 Oracle ICL-9 131 19 0
$P_{1}$ GPT-5-mini Single Store ICL-9 139 11 0
$P_{1}$ GPT-5-mini Shared Store ICL-9 144 6 0
$P_{1}$ GPT-5-mini In Context ICL-9 140 10 0
$P_{1}$ GPT-5-mini Oracle ICL-9 143 7 0
$P_{1}$ GPT-5 Single Store ICL-9 146 4 0
$P_{1}$ GPT-5 Shared Store ICL-9 145 5 0
$P_{1}$ GPT-5 In Context ICL-9 136 14 0
$P_{1}$ GPT-5 Oracle ICL-9 147 3 0
$P_{1}$ GPT-4.1 Single Store N/A 132 18 0
$P_{1}$ GPT-4.1 Shared Store N/A 122 27 1
$P_{1}$ GPT-4.1 In Context N/A 133 17 0
$P_{1}$ GPT-4.1 Oracle N/A 129 21 0
$P_{1}$ GPT-5-mini Single Store N/A 142 8 0
$P_{1}$ GPT-5-mini Shared Store N/A 147 3 0
$P_{1}$ GPT-5-mini In Context N/A 143 7 0
$P_{1}$ GPT-5-mini Oracle N/A 141 9 0
$P_{1}$ GPT-5 Single Store N/A 140 10 0
$P_{1}$ GPT-5 Shared Store N/A 142 8 0
$P_{1}$ GPT-5 In Context N/A 136 14 0
$P_{1}$ GPT-5 Oracle N/A 140 10 0
$P_{2}$ GPT-4.1 Single Store ICL-9 122 28 0
$P_{2}$ GPT-4.1 Shared Store ICL-9 114 35 1
$P_{2}$ GPT-4.1 In Context ICL-9 126 24 0
$P_{2}$ GPT-4.1 Oracle ICL-9 132 18 0
$P_{2}$ GPT-5-mini Single Store ICL-9 145 5 0
$P_{2}$ GPT-5-mini Shared Store ICL-9 133 16 0
$P_{2}$ GPT-5-mini In Context ICL-9 146 4 0
$P_{2}$ GPT-5-mini Oracle ICL-9 142 7 1
$P_{2}$ GPT-5 Single Store ICL-9 139 11 0
$P_{2}$ GPT-5 Shared Store ICL-9 136 12 2
$P_{2}$ GPT-5 In Context ICL-9 138 12 0
$P_{2}$ GPT-5 Oracle ICL-9 142 8 0
$P_{2}$ GPT-4.1 Single Store N/A 130 20 0
$P_{2}$ GPT-4.1 Shared Store N/A 126 23 1
$P_{2}$ GPT-4.1 In Context N/A 126 24 0
$P_{2}$ GPT-4.1 Oracle N/A 140 10 0
$P_{2}$ GPT-5-mini Single Store N/A 145 5 0
$P_{2}$ GPT-5-mini Shared Store N/A 141 9 0
$P_{2}$ GPT-5-mini In Context N/A 140 10 0
$P_{2}$ GPT-5-mini Oracle N/A 141 9 0
$P_{2}$ GPT-5 Single Store N/A 134 15 1
$P_{2}$ GPT-5 Shared Store N/A 144 6 0
$P_{2}$ GPT-5 In Context N/A 140 10 0
$P_{2}$ GPT-5 Oracle N/A 143 6 1
$P_{3}$ GPT-4.1 Single Store ICL-9 128 22 0
$P_{3}$ GPT-4.1 Shared Store ICL-9 120 30 0
$P_{3}$ GPT-4.1 In Context ICL-9 129 21 0
$P_{3}$ GPT-4.1 Oracle ICL-9 129 21 0
$P_{3}$ GPT-5-mini Single Store ICL-9 141 9 0
$P_{3}$ GPT-5-mini Shared Store ICL-9 140 10 0
$P_{3}$ GPT-5-mini In Context ICL-9 142 8 0
$P_{3}$ GPT-5-mini Oracle ICL-9 143 7 0
$P_{3}$ GPT-5 Single Store ICL-9 146 3 1
$P_{3}$ GPT-5 Shared Store ICL-9 147 3 0
$P_{3}$ GPT-5 In Context ICL-9 138 12 0
$P_{3}$ GPT-5 Oracle ICL-9 140 10 0
$P_{3}$ GPT-4.1 Single Store N/A 128 22 0
$P_{3}$ GPT-4.1 Shared Store N/A 121 29 0
$P_{3}$ GPT-4.1 In Context N/A 123 27 0
$P_{3}$ GPT-4.1 Oracle N/A 136 14 0
$P_{3}$ GPT-5-mini Single Store N/A 140 10 0
$P_{3}$ GPT-5-mini Shared Store N/A 140 9 1
$P_{3}$ GPT-5-mini In Context N/A 144 6 0
$P_{3}$ GPT-5-mini Oracle N/A 146 0 0
$P_{3}$ GPT-5 Single Store N/A 133 17 0
$P_{3}$ GPT-5 Shared Store N/A 135 13 2
$P_{3}$ GPT-5 In Context N/A 133 17 0
$P_{3}$ GPT-5 Oracle N/A 139 11 0
$P_{4}$ GPT-4.1 Single Store ICL-9 120 30 0
$P_{4}$ GPT-4.1 Shared Store ICL-9 109 39 2
$P_{4}$ GPT-4.1 In Context ICL-9 125 25 0
$P_{4}$ GPT-4.1 Oracle ICL-9 128 22 0
$P_{4}$ GPT-5-mini Single Store ICL-9 138 12 0
$P_{4}$ GPT-5-mini Shared Store ICL-9 139 11 0
$P_{4}$ GPT-5-mini In Context ICL-9 143 7 0
$P_{4}$ GPT-5-mini Oracle ICL-9 140 10 0
$P_{4}$ GPT-5 Single Store ICL-9 141 9 0
$P_{4}$ GPT-5 Shared Store ICL-9 135 14 1
$P_{4}$ GPT-5 In Context ICL-9 137 12 1
$P_{4}$ GPT-5 Oracle ICL-9 146 4 0
$P_{4}$ GPT-4.1 Single Store N/A 125 25 0
$P_{4}$ GPT-4.1 Shared Store N/A 119 31 0
$P_{4}$ GPT-4.1 In Context N/A 125 25 0
$P_{4}$ GPT-4.1 Oracle N/A 133 17 0
$P_{4}$ GPT-5-mini Single Store N/A 137 12 1
$P_{4}$ GPT-5-mini Shared Store N/A 136 12 2
$P_{4}$ GPT-5-mini In Context N/A 138 12 0
$P_{4}$ GPT-5-mini Oracle N/A 137 13 0
$P_{4}$ GPT-5 Single Store N/A 137 13 0
$P_{4}$ GPT-5 Shared Store N/A 136 13 1
$P_{4}$ GPT-5 In Context N/A 137 13 0
$P_{4}$ GPT-5 Oracle N/A 142 8 0
A.10 Feasibility Analysis
A.10.1 Cost Analysis